OpenStudy (anonymous):

The table below illustrates the decay of a sample of radioactive uranium. Time in Days, x 0 1 2 3 4 5 Sample Remaining (grams), U 500 255 130 66 34 17 Which equation best models this set of data where U represents the amount of sample remaining, in grams, at time x? A. U = 26.7x2 - 223.9x + 481.1 B. U = 500(0.51)x C. U = 500(0.49)x D. U = 392 - 90.2x

OpenStudy (anonymous):

you start with 500 so it is going to look like \(500\times (b)^x\) for some \(b\)

OpenStudy (anonymous):

to find b, take the ratio of any pair of outputs. you get for example \(\frac{255}{500}=.51\) so that is your b

OpenStudy (anonymous):

so what answer choice would that be ?

OpenStudy (anonymous):

@satellite73 what answer choice would that be?

OpenStudy (anonymous):

\(500\times (.51)^x\)

OpenStudy (anonymous):

@satellite73 thank you so much